### v1

# DEGAP: Dual Event-Guided Adaptive Prefixes for Templated-Based Event Argument Extraction Model with Slot Querying

Abstract
Recent advancements in event argument extraction (EAE) involve incorporating beneficial auxiliary information into models during training and inference, such as retrieved instances and event templates. Additionally, some studies introduce learnable prefix vectors to models. These methods face three challenges: (1) insufficient utilization of relevant event instances due to deficiencies in retrieval; (2) neglect of important information provided by relevant event templates; (3) the advantages of prefixes are constrained due to their inability to meet the specific informational needs of EAE. In this work, we propose DEGAP, which addresses the above challenges through two simple yet effective components: (1) dual prefixes, where the instance-oriented prefix and template-oriented prefix are trained to learn information from different event instances and templates, respectively, and then provide relevant information as cues to EAE model without retrieval; (2) event-guided adaptive gating mechanism, which guides the prefixes based on the target event to fully leverage their advantages. Extensive experiments demonstrate that our method achieves new state-of-the-art performance on four datasets (ACE05, RAMS, WIKIEVENTS, and MLEE). Further analysis verifies the importance of the proposed design and the effectiveness of the main components [^1] We will release the code after acceptance..
1 Introduction
As an essential subtask of Event Extraction, Event Argument Extraction (EAE) aims to identify event arguments and their roles for given event mentions in context. For example, in Fig. 1, given a Life.Die event with trigger killed, EAE models need to extract arguments soldiers, bomber, and checkpoint, and the corresponding roles — Victim, Agent, and Place. In the field of NLP, EAE has always been a challenging task, which requires deep natural language understanding to address Hsu et al. (2023).
Consequently, some efforts incorporate ancillary information beneficial for EAE into models during training and inference, such as augmenting input with retrieved instances Du and Ji (2022); Ren et al. (2023); Zhao et al. (2023a) or querying arguments using role slots in event template Liu et al. (2024); Zhang et al. (2022); Ma et al. (2022); He et al. (2023); Nguyen et al. (2023). Moreover, in light of advancements in prompt learning Li and Liang (2021); Liu et al. (2022b), some studies introduce prefixes (i.e, learnable prompt vectors) into models Liu et al. (2022a); Zhang et al. (2023a); Hsu et al. (2023); Nguyen et al. (2023). Despite the empirical success of previous works, three issues remain that have not been appropriately solved, hindering further performance improvements.
Figure 1: Two examples of EAE and event template from Ma et al. (2022). Each underlined part (i.e., role slot) in the template is named after an argument role.
First, relevant event instances are not fully and accurately utilized due to deficiencies in retrieval. Previous studies regard relevant instances retrieved from training data Wang et al. (2022) as cues to extract arguments and learn by analogy. However, due to the need to pre-specify the number of event instances to be retrieved and potentially irrelevant instances retrieved Ren et al. (2023), these methods struggle to fully and accurately utilize relevant reference instances to boost performance. Additionally, these methods often suffer from high computational costs and nontrivial knowledge loss because they independently retrieve and encode event instances, resulting in a poor local optimum Huang et al. (2023b). Here raises an interesting question: Can EAE models benefit from relevant event instances without relying on retrieval?
Second, relevant event templates are not considered. Merely using the corresponding template for the target event as prompt, some important information provided by relevant event templates might be overlooked Ma et al. (2022); He et al. (2023). Moreover, only Ren et al. (2023) attempted to retrieve role labels relevant to the target event as demonstrations and achieved success. As shown in Fig. 1, event templates can provide rich semantic information about events, and relevant events often have similar structures and share partial roles (e.g., Life.Die and Conflict.Attack) Li et al. (2023a). Intuitively, regarding relevant event templates as cues and learning by analogy can help EAE models better extract the target event from complex semantics.
Finally, existing prefixes largely fail to meet the specific informational needs of EAE, leading to unsatisfactory performance. Typically, existing prefixes used for EAE are generalizable across multiple tasks, rather than specifically tailored to a particular task. Despite high portability, recent research indicates that these prefixes do not effectively meet the specific informational needs of the task, failing to fully utilize the advantages of prefix Zhang et al. (2023d). Meanwhile, for the EAE task, some studies show that models should focus on information relevant to the target event and irrelevant information would hinder the model’s performance Xu et al. (2022); Liu et al. (2023c). However, how to adaptively capture relevant information from the prefix at each layer based on the target event, and how to guide the prefix to leverage its advantages fully, remain largely under-explored.
To address these challenges, we propose DEGAP (Dual Event-Guided Adaptive Prefixes for event argument extraction). Inspired by recent studies that utilize prefixes as efficient parameters to learn information from different corpora and integrate retrieved instances into prefixes as cues Zhang et al. (2023b); Liu et al. (2023a), we design two types of prefixes: The instance-oriented prefix is trained to learn event instance information at the event mention-level, and the template-oriented prefix is trained to learn event template information at the ontology level. To adaptively capture relevant information from the prefix at each layer based on the target event and further guide the prefix, we design an event-guided adaptive gating mechanism. It utilizes the target event’s trigger and type to score each instance-oriented and template-oriented prefix token via gated weight assignment at each layer, respectively. Finally, the event-guided prefixes provide relevant information as cues at each layer of the model during inference without retrieval, facilitating the model to extract the target event.
To comprehensively evaluate our method, we conduct extensive experiments on four benchmark datasets (ACE05, RAMS, WIKIEVENTS, MLEE). Experimental results demonstrate that our method outperforms existing methods across all datasets, illustrating the effectiveness of our approach. Furthermore, we conduct a thorough study to explore the reasons for the performance improvement of our method, confirming the significance of the proposed design and the efficacy of main components.
2 Related Work
Event argument extraction. Traditionally, most models are classification-based Chen et al. (2015); Nguyen et al. (2016); Yang et al. (2018); Ebner et al. (2020); Zhang et al. (2020); Du and Cardie (2020a); Xu et al. (2022). Recently, there has been a trend towards framing EAE as a text generation or question-answering (QA) task. Generation-based methods sequentially generate all arguments of the target event by following a pre-defined template Li et al. (2021); Hsu et al. (2021); Zeng et al. (2022); Hsu et al. (2023). QA-based methods query the arguments of each role using corresponding question templates Du and Cardie (2020b); Liu et al. (2021); Wei et al. (2021); Lu et al. (2023); Uddin et al. (2024) or role slots in event template Zhang et al. (2022); Ma et al. (2022); Li et al. (2023a); He et al. (2023).
This work builds on the existing state-of-the-art (SOTA) methods PAIE Ma et al. (2022) and TabEAE He et al. (2023), which solves EAE by using role slots in the event template to query input context for extracting argument spans. With two proposed simple yet effective core components—the dual prefixes and the event-guided adaptive gating mechanism—our method establishes new SOTA performance on four datasets.
Event extraction with auxiliary information. As a task requiring a deep understanding of context, many previous efforts have been put into exploring which auxiliary information is useful for event predictions. One line of studies is incorporating such auxiliary information during training and inference, such as event templates Li et al. (2021); Hsu et al. (2021); Zeng et al. (2022); He et al. (2023); Liu et al. (2024), natural questions Liu et al. (2020, 2021); Lu et al. (2023); Uddin et al. (2024), retrieved instances Du and Ji (2022); Ren et al. (2023); Zhao et al. (2023a), already predicted events Huang et al. (2023a); Du et al. (2022), and syntactic and semantic information Huang et al. (2017); Veyseh et al. (2020); Ahmad et al. (2021); Xu et al. (2022); Yang et al. (2023). However, as mentioned in Section 1, existing methods have not fully exploited the benefits of relevant event instances and event templates.
Prompt learning. Prompt learning is a straightforward and effective method that adjusts PLMs for specific downstream tasks by adding prompts to the input. One line of research applies an automated search to find appropriate discrete prompting words Wallace et al. (2019); Shin et al. (2020); Wen et al. (2024). Another line of research attaches a set of trainable prefix tokens to the model as soft prompts Li and Liang (2021); Liu et al. (2022b); Chen et al. (2023); Li et al. (2023b), which have been widely applied in various studies. Liu et al. (2023a) integrates retrieved instances into the prefix as cues for inference. Zhang et al. (2023b) uses prefixes to learn information from different corpora and refine inference. Additionally, some EAE efforts use prefixes to optimize manually designed prompts Liu et al. (2022a); Nguyen et al. (2023) or to integrate syntactic and semantic information Hsu et al. (2023). Despite the successes, the prefixes used are not specifically tailored for the EAE task, hence the benefits of prefixes can still be further exploited Zhang et al. (2023d).
3 Methodology
In this section, we first define the EAE task. Then, we detail how the proposed DEGAP solves the challenges mentioned.
Task Definition Formally, we first define a set of event types =$\mathcal{E}=\{e_{1},e_{2},e_{3},\ldots\}$ and a corresponding set of roles $R^{e}=\{r_{1},r_{2},r_{3},\ldots\}$ for each event type $e\in\mathcal{E}$. Then, given a context $X=\{x_{1},x_{2},\ldots,x_{n}\}$ of $n$ words, the event trigger $x_{\text{trig}}\in X$ and the event type $e$, the task aims to extract all $(s,r)$ pairs for the event, where $s$ is a text span in $X$, and $r$ is an argument role in $R^{e}$.
3.1 Model Architecture
Figure 2: The overview of DEGAP. Here, the two encoders share the same parameters, and the two decoders also share the same parameters. Due to space constraints, the event template here is simplified. When encoding the context, $\text{Prefix}_{i}$ and $h_{i-1}$ refer to $\text{Prefix}_{i}^{\textit{ins}}$ and $h_{i-1}^{trig}$, respectively. Conversely, they refer to $\text{Prefix}_{i}^{\textit{tem}}$ and $h_{i-1}^{type}$, respectively. IOP: Instance-Oriented Prefix. TOP: Template-Oriented Prefix.
Fig. 2 shows the overall architecture of DEGAP. For given event type $e$, we create an event template for it (§ 3.1.1). To learn and save information from event instances and event templates, we design instance-oriented prefixes and template-oriented prefixes, respectively (§ 3.1.2). Then, we design an event-guided adaptive gating mechanism to adaptively capture relevant information in the prefixes and further guide the prefixes (§ 3.1.3). Subsequently, when encoding the context and event template, these event-guided prefixes provide relevant event information as cues to the model at each layer without retrieval (§ 3.1.4). Finally, we introduce the span selection, which utilizes role slots in the event template to query the context for extracting argument spans (§ 3.1.5).
3.1.1 Event Template Creation
We create an event template $\mathcal{T}_{e}$ for the given event type $e$, utilizing its rich semantic information to guide the model Hsu et al. (2022). Similar to previous works Liu et al. (2022a); Zhang et al. (2023b), the event template $\mathcal{T}_{e}$ consists of two parts: the type part $\mathcal{I}_{e}$ and the schema part $\mathcal{S}_{e}$. The type part $\mathcal{I}_{e}$ adopts the paradigm “Event type is [Event type]”. For the schema part $\mathcal{S}_{e}$, we directly inherit it from TabEAE He et al. (2023). Considering the example in Fig. 2, where there is a Conflict.Attack event with the event trigger hit, the template for this event is: “<s> Event type is Conflict.Attack </s> <s> Attacker (and Attacker) attacked Target (and Target) using Instrument (and Instrument) at Place (and Place) </s>”, where each underlined part (i.e., role slot) is named after an argument role. The brackets indicate that there may be multiple arguments of that role in an event.
3.1.2 Dual Prefixes
To fully benefit from relevant event instances and event templates without retrieval, we construct two types of prefixes: (1) the instance-oriented prefix and (2) the template-oriented prefix. With self-attention mechanisms, these prefixes can learn rich semantic information about events from different instances and templates and further guide the model to extract arguments of the target event.
Instance-Oriented Prefix (IOP). To learn and save information from different event instances, at the $i$-th layer of the encoder or decoder, we equip the self-attention block with a set of learnable prefix tokens oriented towards event instances $\text{Prefix}_{i}^{\textit{ins}}\in\mathbb{R}^{len_{\textit{ins}}\times m}$ , where $len_{\textit{ins}}$ is the length of $\text{Prefix}_{i}^{\textit{ins}}$ and $m$ corresponds to the hidden dimension.
Template-Oriented Prefix (TOP). To learn and save information from different event templates, at the $i$-th layer of the encoder or decoder, we equip the self-attention block [^2] We also attempt to equip the cross-attention block in each layer of the decoder with the prompt-oriented prefix, but observe almost no improvement in performance. with a set of learnable prefix tokens oriented towards event templates $\text{Prefix}_{i}^{\textit{tem}}\in\mathbb{R}^{len_{\textit{tem}}\times m}$, where $len_{\textit{tem}}$ is the length of $\text{Prefix}_{i}^{\textit{tem}}$.
3.1.3 Event-Guided Adaptive Gating mechanism
Not all information within the prefix benefits argument extraction of the target event. To adaptively capture relevant information from the prefix at each layer based on the target event and further guide the prefix, inspired by Liu et al. (2023b), we utilize the target event’s trigger and type to score each prefix token in $\text{Prefix}_{i}^{ins}$ and $\text{Prefix}_{i}^{tem}$ through gated weight assignment, respectively. For example, in the $i$-th layer, the weight distribution of tokens in $\text{Prefix}_{i}^{ins}$ is calculated as:
$a_{i}^{ins}=\text{sigmoid}(h_{i-1}^{trig}W^{ins}_{i}),$ (1)
where $h_{i-1}^{trig}$ is the hidden state of the event trigger from the previous layer obtained through average pooling, and $W^{ins}_{i}\in\mathbb{R}^{m\times len_{ins}}$ corresponds to the parameters to be learned.
Additionally, we also follow Zhang et al. (2023d) to introduce a learnable scaled factor $\lambda_{i}^{ins}$ to balance the information provided by the original input and the prefix. We obtain the final instance-oriented prefix representation $\text{Prefix}_{i}^{ins^{\prime}}$:
$\text{Prefix}_{i}^{ins^{\prime}}=\lambda_{i}^{ins}a_{i}^{ins}\odot\text{Prefix}_{i}^{ins},$ (2)
where $\odot$ is the element-wise multiplication. Similarly, we derive the finally template-oriented prefix representation $\text{Prefix}_{i}^{tem^{\prime}}$:
$a_{i}^{tem}=\text{sigmoid}(h_{i-1}^{type}W^{tem}_{i}),$ (3)
$\text{Prefix}_{i}^{tem^{\prime}}=\lambda_{i}^{tem}a_{i}^{tem}\odot\text{Prefix}_{i}^{tem},$ (4)
where $h_{i-1}^{type}$ is the hidden state obtained through average pooling of the event type from the previous layer, and $W^{tem}_{i}\in\mathbb{R}^{m\times len_{tem}}$ is the learnable parameters.
Finally, the new self-attention module with event information intervention is formalized as:
$H\leftarrow\text{LayerNorm}(H^{{}^{\prime}}+H),$
$H^{{}^{\prime}}=\text{MHSA}(\text{Prefix}_{i}^{{}^{\prime}}\oplus H)_{|\text{Prefix}_{i}^{{}^{\prime}}|:|\text{Prefix}_{i}^{{}^{\prime}}\oplus H|},$ (5)
where $\text{MHSA}(\cdot)$ denotes the multi-head self-attention mechanism, $\text{Prefix}i^{{}^{\prime}}$ represents either $\text{Prefix}{i}^{ins^{\prime}}$ or $\text{Prefix}{i}^{tem^{\prime}}$, $H$ refers to the hidden state, and $(\cdot){b:c}$ denotes the slicing operation on the $seq_{l}en$ dimension from $b$ to $c$.
3.1.4 Encoding with Prefixes
To ensure a fair comparison and to visually demonstrate the effects of our designed prefixes, we adopt the encoding process from prior work He et al. (2023) for encoding context instance and event template, and augment it with our prefixes.
Given input context $X=\{x_{1},x_{2},\ldots,x_{n}\}$ and the event trigger $x_{trig}$ triggering event type $e$, we use text markers <t> and </t> as special tokens and then insert them before and after $x_{trig}$ in $X$, respectively.
$\widetilde{X}=\{x_{1},x_{2},\ldots,\textbf{<t>}x_{trig}\textbf{</t>},\ldots,x_{n}\}.$
Then, we input the processed $\widetilde{X}$ into the encoder and the decoder to obtain the context representation $H_{x}$:
$H_{x}^{enc}=\text{Encoder}(\text{Prefix}_{enc}^{ins},\widetilde{X}),$ (6)
$H_{x}=\text{Decoder}(\text{Prefix}_{dec}^{ins},H_{x}^{enc}),$ (7)
where $\text{Prefix}_{enc}^{ins}$ and $\text{Prefix}_{dec}^{ins}$ are the instance-oriented prefix representations $\text{Prefix}_{i}^{ins}$ from the encoder and decoder, respectively.
For the target event template $\mathcal{T}_{e}$, we input it into the Encoder and the Decoder to obtain the template representation $H_{\mathcal{T}_{e}}$:
$H_{\mathcal{T}_{e}}^{enc}=\text{Encoder}(\text{Prefix}_{enc}^{tem},\mathcal{T}_{e}),$ (8)
$H_{\mathcal{T}_{e}}=\text{Decoder}(\text{Prefix}_{dec}^{tem},H_{\mathcal{T}_{e}}^{enc},H_{x}^{enc}),$ (9)
where $\text{Prefix}_{enc}^{tem}$ and $\text{Prefix}_{dec}^{tem}$ are the template-oriented prefix representations $\text{Prefix}_{i}^{tem}$ from the encoder and decoder, respectively. Note that the cross-attention mechanism is only utilized to decode the event template. When computing Eq. 7, it is skipped.
3.1.5 Span Selection
With the template representation $H_{\mathcal{T}_{e}}$ obtained through Eq. 9, for the $k$-th role slot in the template, we can obtain its representation $h_{s_{k}}$ and span selector $\phi_{k}=\{\phi_{k}^{start},\phi_{k}^{end}\}$ Ma et al. (2022); He et al. (2023):
$\phi_{k}^{start}=h_{s_{k}}\odot w^{start},$
$\phi_{k}^{end}=h_{s_{k}}\odot w^{end},$ (10)
where $w^{start}$ and $w^{end}$ are learnable parameters. Subsequently, the span selector $\phi_{k}$ queries the context to extract at most one argument span $(s_{k},e_{k})$ for role slot $k$:
$\text{logit}_{k}^{start}=\phi_{k}^{start}H_{x}\in\mathbb{R}^{n},$
$\text{logit}_{k}^{end}=\phi_{k}^{end}H_{x}\in\mathbb{R}^{n},$
$\text{score}_{k}(l,r)=\text{logit}_{k}^{start}(l)+\text{logit}_{k}^{end}(r),$
$(s_{k},e_{k})=\operatorname*{arg\,max}_{(l,r)\in\mathcal{C}}\text{score}_{k}(l,r)$ (11)
where candidate span set $\mathcal{C}=\{(l,r)\mid(l,r)\in n^{2},0<r-l<MSL\}\cup\{(0,0)\}$. $MSL$ denotes the max span length.
The loss function is defined as:
$P_{k}^{start}=\text{Softmax}(\text{logit}_{k}^{start}),$
$P_{k}^{end}=\text{Softmax}(\text{logit}_{k}^{end}),$
$\mathcal{L}=-\sum_{X}\sum_{k}[\log P_{k}^{start}(s_{k})+\log P_{k}^{end}(e_{k})],$ (12)
where $X$ ranges over all context in dataset and $k$ ranges over all role slots in the template for $X$, and $(s_{k},e_{k})$ is the golden span for role slot $k$. For a role slot that does not relate to any argument, it is assigned the empty span $(0,0)$.
4 Experiments
4.1 Experimental Setup
Model PLM ACE05 RAMS WIKIEVENTS MLEE
Arg-I Arg-C Arg-I Arg-C Arg-I Arg-C Arg-I Arg-C
TSAR Xu et al. (2022) RoBERTa - - 57.0 52.1 71.1 65.8 72.6 71.5
SCPGR Liu et al. (2023c) RoBERTa - - 56.5 52.2 71.5 66.2 72.5 71.3
BART-Gen Li et al. (2021) RoBERTa 69.9 66.7 51.2 47.1 66.8 62.4 71.0 69.8
DEGREE Hsu et al. (2022) BART 75.6 73.0 - - - - - -
AMPERE Hsu et al. (2023) BART 76.0 73.4 - - - - - -
APE Zhang et al. (2023b) BART 75.3 72.9 56.3 51.7 70.6 65.8 - -
R-GQA Du and Ji (2022) BART 75.6 72.8 - - - - - -
AHR Ren et al. (2023) T5 - - 54.6 48.4 69.6 63.4 - -
PAIE Ma et al. (2022) BART 75.7 72.7 56.8 52.2 70.5 65.3 72.1 70.8
SPEAE Nguyen et al. (2023) BART - - 58.0 53.3 71.9 66.1 - -
TabEAE He et al. (2023) RoBERTa 75.5 72.6 57.0 52.5 70.8 65.4 71.9 71.0
DRAAE (Ours) RoBERTa 76.1 73.2 57.5 53.0 71.0 66.0 72.8 71.7
DEGAP (Ours) RoBERTa 76.6 74.4 58.5 54.2 72.2 67.1 74.0 73.4
Table 1: The overall performance of all methods. Models in the PLM column are of large-scale (with 24 Transformer layers). The highest scores are highlighted in bold, and the second-highest scores are underlined.
Datasets. We conduct experiments on four benchmark datasets, specifically including ACE05 Doddington et al. (2004), RAMS Ebner et al. (2020), WIKIEVENTS Li et al. (2021), and MLEE Pyysalo et al. (2012). ACE05 is a sentence-level dataset, while the other three are document-level datasets. We provide detailed introductions to these datasets in Appendix A.
Evaluation Metrics. We adopt two evaluation metrics: (1) Argument Identification F1 score (Arg-I): A predicted argument for an event is correctly identified if its boundary matches any golden arguments of the event. (2) Argument Classification F1 score (Arg-C): a predicted argument of an event is correctly classified only if its boundary and role type both match the ground truth. The final reported performance is averaged over five random runs.
Implementation details. Please refer to Appendix B.1 for details of implementation.
Baselines. Details about the baselines are listed in Appendix B.2. For a comprehensive comparison, we also conduct a comparative analysis between DEGAP and LLMs in Appendix C.1.
We also compare our method with its variant, the Dual Retrieval-Augmented Argument Extraction model (DRAAE), which retrieves relevant event instances and templates as cues to guide the model, instead of using the prefixes we designed. We provide further implementation details of this model in Appendix B.3.
4.2 Overall Performance
Table 1 presents the performance of all baselines and our model across all datasets. From the results, we can conclude:
(1) Our method significantly outperforms all the baselines and achieves new SOTA results on the four datasets. Specifically, our method improves the Arg-C F1 score by 1.0%, 0.9%, 0.9%, and 1.9% on ACE05, RAMS, WIKIEVENTS, and MLEE, respectively. This demonstrates the efficacy of our method in handling both sentence-level EAE and document-level EAE tasks.
(2) Compared to methods using retrieval (i.e., R-GQA and AHR) and those using event templates (e.g., DEGREE and TabEAE), our approach shows better performance on all datasets. We attribute this to DEGAP’s ability to fully utilize relevant event instance and event template information as cues. In contrast, previous methods provide insufficient guidance to the model from only one perspective.
(3) Compared to methods that use prefixes, such as AMPERE and SPEAE, which utilize auxiliary semantic information to augment their prefixes, our method does not rely on these modules and still surpasses them to varying degrees. This suggests that our event-guided prefixes are more suitable for EAE, thereby leveraging their advantages.
(4) Our DRAAE surpasses the baseline model TabEAE across all datasets. This indicates that augmenting input with relevant event instances and event templates is effective for extracting arguments. However, our DEGAP still outperforms DRAAE, primarily because it addresses two shortcomings in retrieval: a) the need to pre-specify the number of relevant instances and templates to be retrieved, and potentially irrelevant retrieval results, leading to insufficient utilization of relevant semantic information; b) the utilization of pre-trained retriever with frozen parameters, leading to less relevance to the EAE task and yielding a poor local optimum. Moreover, the efficiency analysis in Appendix C.2 indicates that due to our DEGAP not requiring retrieval, it has a faster inference speed.
Model ACE05 ACE05 ACE05 ACE05 RAMS RAMS RAMS RAMS WIKI WIKI WIKI WIKI MLEE MLEE MLEE MLEE Avg
$\Downarrow$ $\Downarrow$ $\Downarrow$ $\Downarrow$ $\Downarrow$ $\Downarrow$ $\Downarrow$ $\Downarrow$ $\Downarrow$ $\Downarrow$ $\Downarrow$ $\Downarrow$ $\Downarrow$ $\Downarrow$ $\Downarrow$ $\Downarrow$
ACE05 RAMS WIKI MLEE RAMS ACE05 WIKI MLEE WIKI ACE05 RAMS MLEE MLEE ACE05 RAMS WIKI
TabEAE 72.6 20.1 41.2 5.4 52.5 32.3 20.1 19.6 65.4 48.7 32.6 9.6 71.0 8.5 14.4 9.9 32.7
DEGAP 74.4 23.8 44.8 11.5 54.2 33.9 27.6 25.7 67.1 50.1 34.4 29.3 73.4 21.9 19.1 22.5 38.4
[…]
Model ACE05 RAMS WIKI MLEE
S${}_{[\textbf{185}]}$ M${}_{[\textbf{218}]}$ S${}_{\textbf{[587]}}$ M${}_{\textbf{[284]}}$ S${}_{\textbf{[114]}}$ M${}_{\textbf{[251]}}$ S${}_{\textbf{[175]}}$ M${}_{\textbf{[2025]}}$
TabEAE 71.2 73.8 52.8 51.6 65.3 65.5 79.3 70.3
DEGAP 72.7 (1.5) 75.8 (2.0) 54.4 (1.6) 54.0 (2.4) 66.5 (1.2) 67.3 (1.8) 81.1 (1.8) 72.8 (2.5)
Table 3: Performance (Arg-C F1 score) on instances with different numbers of events in context. The numbers in square brackets indicate the number of supporting events in the test set. S: Single event. M: Multiple events.
Model ACE05 RAMS WIKI MLEE
N-O${}_{\textbf{[319]}}$ Overlap ${}_{\textbf{[84]}}$ N-O${}_{\textbf{[690]}}$ Overlap${}_{\textbf{[181]}}$ N-O${}_{\textbf{[296]}}$ Overlap${}_{\textbf{[69]}}$ N-O${}_{\textbf{[1466]}}$ Overlap${}_{\textbf{[734]}}$
TabEAE 71.1 78.6 51.6 55.6 65.7 64.4 75.4 65.8
DEGAP 72.1 (1.0) 80.5 (1.9) 52.9 (1.3) 59.5 (3.9) 67.3 (1.6) 66.7 (2.3) 76.9 (1.5) 68.5 (2.7)
Table 4: Results (Arg-C F1 score) of extracting overlapping (events with shared arguments) and non-overlapping events. Numbers in square brackets indicate the number of supporting events in the test set. N-O: Non-overlapping.
5 Analysis
In this section, we conduct a comprehensive study to analyze the reasons for the performance improvement of our method, and to validate the significance of our proposed design and the effectiveness of the main components.
5.1 Why is DEGAP Effective?
We explore from two perspectives: a) domain transfer and b) handling complex contextual semantics, to investigate whether the performance gains in our method are due to DEGAP providing the model with relevant event instances and event templates as cues. We consider the base model TabEAE.
Domain Transfer. We examine the performance of the model when facing domain transfer. Specifically, we select four different domain datasets: ACE05, RAMS, WIKIEVENTS, and MLEE, train the model on one domain dataset (src), and then test it on all domain datasets (tgt). The results are shown in Table 2. Compared to the results w/o domain transfer, both our method DEGAP and the baseline TabEAE exhibit a sharp decline under different src-tgt transfer settings, where the baseline’s performance even falls below 10% in some cases (e.g., WIKI $\Rightarrow$ MLEE). This indicates that extracting arguments from new types of events is highly challenging for the model. However, encouragingly, our method DEGAP consistently outperforms the baseline in different src-tgt transfer settings. We attribute this to DEGAP’s ability to use relevant semantic information in the source domain as cues and learn by analogy to better understand new types of events in the target domain.
Handing Complex Contextual Semantics. We explore the ability of models to handle complex semantics from two perspectives: the number of events and the degree of event relevance. (1) Number of events. Multiple events in the same context often exhibit complex semantic connections. We measured the performance of the models on instances whose context contains different numbers of events. As shown in Table 3, compared to the baseline TabEAE, our method shows significant improvements in performance across all datasets, especially in instances that contain multiple events within the context. (2) Degree of event relevance. If two events in the context share arguments, their semantic connection is often more complex. We measured the performance of the model in extracting arguments from overlapping events (events with shared arguments). Table 4 shows that our method achieves the best performance across all datasets, particularly gaining greater improvements in handling overlapping events.
Experimental results support our conclusion: by fully utilizing relevant event instances and event templates to guide the model, our method excels in identifying the target event from complex contextual semantics.
5.2 Why are Dual Prefixes Necessary?
No. Model ACE05 RAMS WIKI MLEE
1 DEGAP 74.4 54.2 67.1 73.4
2 DEGAP-CCT 73.6 53.7 66.6 72.3
3 DEGAP-SP 73.9 53.4 66.8 72.6
4 only IOP (w/o TOP) 73.8 52.9 66.4 72.6
5 only TOP( w/o IOP) 73.6 52.7 66.3 72.7
6 w/o DEGAP 72.8 52.0 65.7 71.2
7 DEGAP-TST 73.9 53.0 66.6 72.8
Table 5: Ablations (Arg-C F1 score) on dual prefixes. Note that all models here are equipped with our proposed event-guided adaptive gating mechanism.
A core question is why we need to design separate prefixes for event instances and event templates. To explore whether the design of dual prefixes is necessary in our method, we compare DEGAP with two single-prefix variants:
  • DEGAP-CCT (Concatenate Context and Template): Instead of encoding the context and template separately, this variant directly concatenates them and then inputs the combined sequence into the Encoder, which is equipped with a set of shared prefixes. For a fair comparison, we use RoBERTa-large as the encoder.
  • DEGAP-SP (Shared Prefix): This variant follows the same encoding process as DEGAP but replaces the instance-oriented and template-oriented prefixes with a single set of shared prefixes.
Moving forward, we further investigate the effects of instance-oriented prefix (IOP) and template-oriented prefix (TOP). Finally, we consider DEGAP-TST (event-Template-Specific TOP), where each event template has its specific TOP. Therefore, TOP solely learns and provides semantic information specific to the target event template, without considering other relevant event templates. Table 5 shows the results.
From the table, we observe that compared to our DEGAP (line 1), the single-prefix variants DEGAP-CCT and DEGAP-SP (lines 2-3) show performance declines to varying degrees across all datasets. We speculate that this may be due to the prefixes becoming confused by simultaneously learning information about both event templates and event instances. Therefore, we choose to use dual prefixes instead of single prefixes in the proposed method.
Additionally, compared to the baseline w/o DEGAP (line 6), introducing IOP or TOP separately will result in a performance increase on all datasets (lines 4-5), and the performance gains by introducing them at the same time exceed the performance improvement caused by introducing any one module alone (line 1). This indicates that it is necessary to utilize IOP and TOP to respectively provide the model with relevant event instance information and event template information as cues for inference. Moreover, IOP and TOP can offer complementary guidance to the model from the event mention-level and ontology level, respectively.
Finally, an interesting phenomenon is that despite DEGAP-TST introducing more parameters than DEGAP, its performance declines across all four datasets (line 7). This further confirms the necessity of providing the model with relevant event template information as cues.
5.3 Is Event-Guided Adaptive Gating mechanism Necessary?
No. Model ACE05 RAMS WIKI MLEE
1 DEGAP 74.4 54.2 67.1 73.4
2 w/o EGAG 73.4 53.3 66.4 72.3
3 DEGAP-S 73.9 53.6 66.8 72.6
Table 6: The study (Arg-C F1 score) of event-guided adaptive gating mechanism.
To explore the importance of the Event-Guided Gating Adaptive mechanism (EGAG), we compare DEGAP with the following variants:
  • w/o EGAG: To investigate whether the performance improvement of DEGAP is due to the additional parameters introduced (gating weights and scaling factors) or because the event-guided prefixes are truly more suitable for EAE, we designed the variant DEGAP w/o EGAG. For a fair comparison, we expanded the prefix length to 1.5 times to align its parameter count with DEGAP.
  • DEGAP-S: To validate the importance of using the target event’s trigger and type to guide IOP and TOP respectively, we follow Zhang et al. (2023d) to directly use the hidden state of the first input token “<s>” in the previous layer as $h_{i-1}^{ins}$ in Eq. 1 and $h_{i-1}^{tem}$ in Eq. 4, respectively.
Table 6 shows that removing EGAG results in a noticeable decline in Arg-C F1 scores across the four datasets. This phenomenon indicates that event-guided prefixes are indeed more suitable for EAE. Additionally, DEGAP-S also shows a decrease in Arg-C F1 scores across the four datasets. This may be because it uses only the first input token “<s>” to calculate $h_{i-1}^{ins}$ and $h_{i-1}^{tem}$, which is insufficient to capture information relevant to the target event in the prefixes. However, compared to the model w/o EGAG, DEGAP-S shows improved performance across all four datasets. This further confirms that previous prefixes do not fully meet the specific informational needs of EAE and leverage their advantages.
We further demonstrate the effects of EGAG visually on the ACE05 dataset.
Attention Visualization. We randomly select a sample and visualize the weight distribution of prefix tokens in IOP and TOP. As shown in Fig. LABEL:prefix_token_weight, the target event has different levels of attention to the prefixes of different layers. More specifically, even within the same layer, the attention to different prefix tokens varies. This observation confirms that EGAG can adaptively capture information from the prefix tokens at each layer based on the target event and further guide the prefixes.
Prefix length. We further analyze the impact of prefix length on model performance. Intuitively, a longer prefix implies more trainable parameters, thus enabling it to provide more expressive guidance for the model regarding relevant semantic information. As shown in Fig. LABEL:prefix_len, as the prefix length increases, the model performance tends to improve, followed by a slight decline. We speculate that this is because overly long prefixes provide redundant semantic information, making it difficult for the model to focus on extracting the target event.
6 Conclusion
In this work, we introduce DEGAP. Based on observations of existing research and comprehensive experimental analysis, we show that prefixes can effectively provide relevant event information to models as cues. Additionally, we introduce an event-guided adaptive gating mechanism that guides prefixes to adapt to different events to fully leverage their advantages. Through extensive experiments, we demonstrate the validity of our newly introduced methods. In the future, we plan to integrate more beneficial auxiliary information into prefixes for solving EAE.
Limitations
We summarize the limitations of our method as follows:
  • Although experimental results indicate that our method exhibits profound performance, our designed prefixes lack sufficient interpretability compared to manually crafted prompts.
  • Due to limited computational resources, we only equipped our prefixes for use on PLMs. In fact, some other works have successfully integrated these efficient prefix parameters into LLMs (e.g., Llama 2 Touvron et al. (2023)) and demonstrated excellent results Zhang et al. (2023c). How to effectively combine our prefixes with these LLMs for solving the EAE task remains an open research question.
  • Compared to other methods, our approach requires additional time costs to find suitable prefix lengths for both IOP and TOP.
  • While our approach is immune to high computational costs and inefficient use of information from relevant instances and templates due to not requiring retrieval, it also presents a significant challenge in how to benefit from other useful semantic information beyond the training corpus. In contrast, retrieval-based methods can leverage richer external semantic information by expanding the retrieval source.
We will attempt to address these issues in future work.
Appendix A Datasets
Dataset ACE05 RAMS WIKIEVENTS MLEE
# Event types 33 139 50 23
# Args per event 1.19 2.33 1.40 1.29
# Events per text 1.35 1.25 1.78 3.32
# Events
Train 4202 7329 3241 4442
Dev 450 924 345 -
Test 403 871 365 2200
Table 7: Statistics of datasets.
ACE05 Doddington et al. (2004) [^3] https://catalog.ldc.upenn.edu/LDC2006T06 is a joint information extraction dataset that compiles news, telephone conversations, and broadcast news, providing annotations for entities, relations, and events in English, Chinese, and Arabic. We utilize only its English event annotations for sentence-level EAE task, adopting the data preprocessing methods used by Wadden et al. (2019). This dataset assumes that events are contained within a single sentence, rather than in a document, which is more typical in real-world scenarios.
RAMS Ebner et al. (2020) [^4] https://nlp.jhu.edu/rams/ is a document-level EAE dataset focused on English online news. Each example is a document consisting of five sentences, where the trigger word indicates a pre-defined event type, and their arguments are dispersed throughout the document. Therefore, a major challenge of this dataset is to address the long-distance dependencies between event triggers and arguments.
WIKIEVENTS Li et al. (2021) [^5] https://github.com/raspberryice/gen-arg is a document-level EAE dataset. The documents in this dataset are derived from English Wikipedia articles that describe real-world events, followed by the retrieval of related news articles based on reference links. They also annotate coreference links for arguments; however, in this task, we only utilize annotations for conventional arguments. Compared to RAMS, WIKIEVENTS contains a greater number of multi-event documents.
MLEE Pyysalo et al. (2012)666 http://www.nactem.ac.uk/MLEE/ is a document-level event extraction dataset containing manually annotated summaries of biomedical publications written in English. We follow the preprocessing procedure outlined by Trieu et al. (2020). As the preprocessed data only includes training/test splits, we follow He et al. (2023) to use the training set as the development set. Moreover, overlapping events are more common in MLEE than in the other three datasets.
Data Statistics. Table 7 shows detailed statistics for each dataset.
Hyperparameters Search Range ACE05 RAMS WIKIEVENTS MLEE
Batch Size [2, 16] 16 4 4 4
Training Steps - 10000 10000 10000 10000
Window Size - 250 250 250 250
Max Span Length - 10 10 10 10
Learning Rate - 2e-5 2e-5 2e-5 2e-5
Warmup Ratio - 0.1 0.1 0.1 0.1
Max Gradient Norm - 5 5 5 5
Max Encoder Seq Length - 200 500 500 500
Max Decoder Seq Length - 250 200 360 360
IOP Length [10, 100] 70 40 40 40
TOP Length [5, 50] 30 20 10 10
Table 8: Hyperparameters for DEGAP. The hyperparameters without a search range are set to the same values as in He et al. (2023).
Appendix B Experimental Setup
B.1 Implementation details
We implement DEGAP using PyTorch and conduct experiments on a single NVIDIA GeForce RTX 3090 24G. For simplicity, we randomly initialize the embeddings of the prefix. Following TabEAE’s setting He et al. (2023), we use the first 17 layers of RoBERTa-large Liu et al. (2019) to instantiate the encoder, and the weights of the remaining 7 layers to initialize the weight of self-attention layers and feedforward layers of the decoder. We also randomly initialize the weight of cross-attention parts of the decoder. The optimization of our model employs the AdamW optimizer Loshchilov and Hutter (2017) equipped with a linear learning rate scheduler. Empirical studies on the selection of the number of encoder/decoder layers can be found in He et al. (2023).
Most hyperparameters are the same as those in TabEAE He et al. (2023). For hyperparameters requiring tuning, we conduct grid searches in specific ranges for each dataset, choosing values that maximize the Arg-C F1 on the development set. The search ranges and the final hyperparameter values are listed in Table 8.
B.2 Baselines
We compare our method with the following state-of-the-art models:
  • TSAR Xu et al. (2022): A classification-based model that first identifies candidate spans and then predicts their roles within events. It utilizes an abstract meaning representation (AMR) graph to help learn the better representations of candidate spans. We report the results from He et al. (2023).
  • SCPRG Liu et al. (2023c): A classification-based model that enhances the final candidate span representations using context information filtered by event trigger and role interaction features. We use their code [^7] https://github.com/LWL-cpu/SCPRG-master to test its performance on all datasets.
  • Bart-Gen Li et al. (2021): A conditional generation model that uses a sequence-to-sequence model and pre-defined templates to sequentially generate arguments. We report the results from He et al. (2023).
  • DEGREE Hsu et al. (2022): A conditional generation model. In addition to templates, event keywords, and event descriptions are also used as auxiliary information. We report the results from Hsu et al. (2023).
  • AMPERE (Hsu et al., 2023): A model that builds upon DEGREE by integrating AMR information into the model using prefixes. Although comparing it directly with our method may not be fair (as it additionally trains an AMR encoder implemented by BART-large or RoBERTa-large, thus has about twice the parameters of our model), we still list the performance of its BART version for reference.
  • APE Zhang et al. (2023b): A conditional generation model. The original paper offers Multi and Single versions trained on multiple and single datasets, respectively. For a fair comparison, we report the performance of its Single version.
  • AHR Ren et al. (2023): A model that utilizes retrieved relevant instances to facilitate sequence generation. The original paper offers three retrieval strategies: Context-Consistency Retrieval, Schema-Consistency Retrieval, and Adaptive Hybrid Retrieval. We report their best-performing method, Adaptive Hybrid Retrieval (AHR).
  • R-GQA Du and Ji (2022): A model that generates arguments given a retrieved demonstration and natural questions. We report the results from the original paper.
  • PAIE Ma et al. (2022): A QA-based model that extracts arguments by querying context using role slots in the event template. We report the results from He et al. (2023).
  • SPEAE Nguyen et al. (2023): A model based on PAIE that utilizes prefixes to integrate relevant context information into the model and alleviate sub-optimal manual prompts. We report the results from the original paper.
  • TabEAE He et al. (2023): Building upon PAIE, it simultaneously extracts multiple events and replaces the base model BART used by PAIE with RoBERTa. The original paper introduces three training-inference paradigms: Multi-Multi, Multi-Single, and Single-Single (consistent with our training-inference paradigm). To ensure a fair comparison, we report the performance of its Single-Single version.
B.3 The implementation details of DRAAE
We follow Ren et al. (2023) to use a frozen S-BERT Reimers and Gurevych (2019) to retrieve the top-$k_{ins}$ relevant instance $\textbf{d}_{ins}$ and the top-$k_{tem}$ relevant templates $\textbf{d}_{tem}$ from training corpus for the input instance and template, respectively. Then, we use RoBERTa-large as an encoder to obtain the retrieved instance embedding $\textbf{h}_{\textbf{d}_{ins}}$ and template embedding $\textbf{h}_{\textbf{d}_{tem}}$ through the representation of the first input token:
$\textbf{h}_{\textbf{d}_{ins}}=\text{RoBERTa}(\textbf{d}_{ins}),$
$\textbf{h}_{\textbf{d}_{tem}}=\text{RoBERTa}(\textbf{d}_{tem}),$ (13)
where $\textbf{h}_{\textbf{d}_{ins}}=\{\textbf{h}_{\textbf{d}_{ins}}^{(1)},\textbf{h}_{\textbf{d}_{ins}}^{(2)},...,\textbf{h}_{\textbf{d}_{ins}}^{(k_{ins})}\}$, and $\textbf{h}_{\textbf{d}_{tem}}=\{\textbf{h}_{\textbf{d}_{tem}}^{(1)},\textbf{h}_{\textbf{d}_{tem}}^{(2)},...,\textbf{h}_{\textbf{d}_{tem}}^{(k_{tem})}\}$. These retrieved instance contexts and event templates are incorporated into each transformer layer’s self-attention module.
The new self-attention with retrieved event information is formulated as:
$H\leftarrow\text{LayerNorm}(H^{{}^{\prime}}+H),$
$H^{{}^{\prime}}=\text{MHSA}(\textbf{h}_{\textbf{d}}\oplus H)_{|\textbf{h}_{\textbf{d}}|:|\textbf{h}_{\textbf{d}}\oplus H|},$ (14)
where $\textbf{h}_{\textbf{d}}$ represents either $\textbf{h}_{\textbf{d}_{ins}}$ or $\textbf{h}_{\textbf{d}_{tem}}$, $H$ is the hidden state.
Finally, we use RoBERTa in Eq. 13 to encode the input context and event template in the same manner as in DEGAP (§ 3.1.4), and further extract argument spans (§ 3.1.5).
Appendix C Experimental Analysis
C.1 Performance Comparison with LLMs
In this section, we compare our method with three publicly available LLMs: GPT-3 Brown et al. (2020) in its text-davinci-003 and GPT-3.5-turbo-instruct versions Ouyang et al. (2022), and GPT-4 Achiam et al. (2023).
Model RAMS
Arg-I Arg-C
text-davinci-003 Brown et al. (2020) 46.2 39.6
GPT-3.5-turbo-instruct Ouyang et al. (2022) 43.3 37.0
GPT-4 Achiam et al. (2023) 52.4 44.1
EDGAP (Ours) 58.5 54.2
Table 9: Performance comparison with LLMs. We report their performance from Zhou et al. (2023).
Table 9 shows that our method achieves improvements of 6.1% and 10.1% in Arg-I and Arg-C F1 scores respectively, emphasizing our method’s significant advantage over LLMs. This makes sense because our model are trained using task-specific data, whereas LLMs can only perform in-context learning through one or two demonstrations. This observation is also consistent with previous findings, suggesting that despite LLMs making substantial progress in many natural language processing tasks Zhao et al. (2023b); Muennighoff et al. (2022) such as text generation tasks Qin et al. (2023), further efforts are still needed for them to excel in generating complex structured outputs (e.g., generating structured event information) Tang et al. (2023); Chen et al. (2024).
C.2 Efficiency Analysis
Model ACE05 RAMS WIKIEVENTS MLEE
DRAAE 61.18 256.87 121.76 626.31
TabEAE 3.69 41.35 18.11 71.96
DEGAP 4.36 43.36 19.64 88.60
Table 10: Inference time (second) for different models on test set of four datasets. Experiments are run on one same NVIDIA GeForce RTX 3090.
Table 10 reports the overall inference time for DEGAP (w/o retrieval) and DRAAE (w/ retrieval). It can be observed that DEGAP runs approximately 6-7 times faster than DRAAE on RAMS, WIKIEVENTS, and MLEE, and approximately 15 times faster than DRAAE on ACE05. This is mainly because: 1) DRAAE requires additional time costs to retrieve the top-$k$ similar event instances and templates; 2) For the retrieved event instances and templates, DRAAE needs to feed them into the encoder to obtain meaningful embeddings. In contrast, DEGAP does not require retrieval and further encoding. Only equipped with a set of learnable prefixes, our DEGAP can adaptively learn semantic information from different event instances and templates during training, and provide relevant guidance at each model layer during inference. Note that due to the parallelized attention calculation for the introduced prefixes and original input on the GPU, there is a negligible impact on inference speed compared to the base model TabEAE.
C.3 Event Template Variant
As described in § 3.1.1, the event template $\mathcal{T}_{e}$ consists of two parts: the type part $\mathcal{I}_{e}$ and the schema part $\mathcal{S}_{e}$. To explore the necessity of type part $\mathcal{I}_{e}$, following Shi et al. (2023), we use type markers “<Event type>” and “</Event type>” as learnable special tokens, inserting them respectively into the corresponding schema part $\mathcal{S}_{e}$ as the template for event type $e$. Additionally, we compute Eq. 3 by averaging the representations of the type markers.
Model ACE05 RAMS WIKIEVENTS MLEE
w/ type part 74.4 54.2 67.1 73.4
r/ type part 74.2 53.9 67.3 73.1
Table 11: Analysis of the necessity for the Type part. The numbers in the table all refer to the Arg-C F1 scores. r/: Replace the type part with type markers.
The experimental results in Table 11 suggest that describing the event type with natural language sentence generally outperforms the use of type markers. Consequently, we retain this setting in our event template $\mathcal{T}_{e}$. Note that as our method does not rely on type markers that need to be learned, it can more easily adapt to new event types.